OpenStudy (anonymous):

What is the slope of the line passing through (5,-1) and (5,3)? A) Undefined B) 1/5 C) 0 D) 5

OpenStudy (anonymous):

slope bw two points (x1,y1)and (x2,y2) can be find by (y2-y1)/(x2-x1)

OpenStudy (anonymous):

I got 4

OpenStudy (anonymous):

3-(-1)/5-5=4

OpenStudy (anonymous):

no its infinite because\[4/0= \infty\]

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

so slope is 90 degree

OpenStudy (anonymous):

That's not any of the answers though

OpenStudy (cggurumanjunath):

4/0 = infinity !

OpenStudy (tkhunny):

The slope is NOT 90º. The line makes a 90º angle with the x-axis. The denominator being zero suggests that the "slope" does not exist. That IS in the given answers.

OpenStudy (cggurumanjunath):

undefined !

OpenStudy (cggurumanjunath):

infinity or undefined.

OpenStudy (anonymous):

So is it 0 or undefined

OpenStudy (anonymous):

oh sorry.

OpenStudy (cggurumanjunath):

undefined.

OpenStudy (tkhunny):

Asked and answered. You tell me if it's 0 or undefined based on the fact that it doesn't exist. Zero exists!

OpenStudy (anonymous):

Oh, lol that makes sense

OpenStudy (anonymous):

Thanks to everyone!